find doman: I don't understand how to find it.
12/(x^2-5x)

Answer by lwsshak3(11628)

find doman: I don't understand how to find it.
12/(x^2-5x)
To find the domain of a function, you must identify value(s) of x that makes the function invalid.
In this case, if the denominator=0, the function becomes undefined.
So set the the denominator=0 and solve for x.
x^2-5x=0
x(x-5)=0
..
x≠0
or
x-5=0
x≠5
Domain: all real numbers except 0 and 5
(-∞,0) U (0,5) U (5,∞)